A scientist has 200 grams of a substance which has half-life of 5yrs. The graph of its decay function is represented by the equation y=ab^x where x is the number of 15 yr periods. What are the values of a and b
Answer by stanbon(75887)

A scientist has 200 grams of a substance which has half-life of 5yrs. The graph of its decay function is represented by the equation y=ab^x where x is the number of 15 yr periods. What are the values of a and b
-------
y = ab^x
-----
y = 200(1/2)^(t/5)
----
y = 200[(1/2)^(1/5)]^t
---
a = 200
b = (1/2)^(1/5)
=====================
Cheers,
Stan H.